OpenStudy (anonymous):

College Algebra: I'm reviewing for my Final coming up and I need tons of help. Some of the math problems I'm going through are not making sense. The math problem will be in the comments:

OpenStudy (anonymous):

\[\sqrt{x+10}=x-2\]

jimthompson5910 (jim_thompson5910):

You have the square root isolated on the left side. What you can do is square both sides to undo the square root (which effectively gets rid of the square root).

jimthompson5910 (jim_thompson5910):

What do you get when you do this?

OpenStudy (anonymous):

Wouldn't you try to get x alone on one side?

jimthompson5910 (jim_thompson5910):

Yes that is the ultimate goal, but first you need to get rid of the square root.

OpenStudy (anonymous):

I thought we got rid of the square root

jimthompson5910 (jim_thompson5910):

Well you have \(\sqrt{x+10}=x-2\) right?

jimthompson5910 (jim_thompson5910):

I see the square root on the left side.

OpenStudy (anonymous):

Yes. You said we square root both sides, so wouldn't it be: \[\sqrt{x+10}=\sqrt{x-2}\]

jimthompson5910 (jim_thompson5910):

Square, not square root

OpenStudy (anonymous):

oops lolx

jimthompson5910 (jim_thompson5910):

Try it again.

OpenStudy (anonymous):

So it's more like \[(x+10=x-2)^{2}\] or \[x+10=(x-2)^{2}\]

jimthompson5910 (jim_thompson5910):

The notation would look like this: \[\Large \sqrt{x+10}=x-2\] \[\Large (\sqrt{x+10})^2=(x-2)^2\] \[\Large x+10=(x-2)^2\] Notice how squaring the \(\sqrt{x+10}\) makes it turn into \(x+10\)

OpenStudy (anonymous):

So, I was right; yay! :)

jimthompson5910 (jim_thompson5910):

yes be sure to apply the square to both sides where each side gets their own square exponent

OpenStudy (anonymous):

Okay, so the equation would be x+10=x^2 +4?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

\((x-2)^2\) is NOT equal to \(x^2+4\) (common mistake)

OpenStudy (anonymous):

Oh?

jimthompson5910 (jim_thompson5910):

You have to use the FOIL rule.

jimthompson5910 (jim_thompson5910):

\((x-2)^2 = (x-2)(x-2) = ??\)

OpenStudy (anonymous):

Oh I got it. \[x ^{2}-4x+4\]

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

You now should have \[\Large x+10 = x^2-4x+4\]

OpenStudy (anonymous):

Okay, i do

jimthompson5910 (jim_thompson5910):

Do you see how to solve from here?

OpenStudy (anonymous):

I'm thinking of moving what's on the left to the right

jimthompson5910 (jim_thompson5910):

That will work.

OpenStudy (anonymous):

Alright, I got \[x ^{2}-5x-6\]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's equal to 0

OpenStudy (anonymous):

Oops, I mean \[x ^{2}-5x-6=0\]

jimthompson5910 (jim_thompson5910):

That's ok, I knew what you meant.

OpenStudy (anonymous):

Okay good :D

OpenStudy (anonymous):

So, that's the answer?

jimthompson5910 (jim_thompson5910):

No, you need to isolate x. You can factor, complete the square or use the quadratic formula.

OpenStudy (anonymous):

I'll use the quadratic formula

jimthompson5910 (jim_thompson5910):

That's a good idea since the quadratic formula works for any quadratic equation.

OpenStudy (anonymous):

Yeah, plus it's my favorite formula :D

OpenStudy (anonymous):

Alright, as far as I got. It's \[x=\frac{ 5\pm7 }{ 2 }\]

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so what are the two possible solutions?

OpenStudy (anonymous):

\[x=\frac{ 5+7 }{ 2 }\] and \[x=\frac{ 5-7 }{ 2 }\]

OpenStudy (anonymous):

wait a minute...

jimthompson5910 (jim_thompson5910):

you can keep going to simplify

OpenStudy (anonymous):

Yeah, that's what I was thinking. So the two solutions are 6 and -1

jimthompson5910 (jim_thompson5910):

They are the possible solutions, so you'll need to check them both in the original equation.

OpenStudy (anonymous):

Only 6 works because -1 makes the equation false

jimthompson5910 (jim_thompson5910):

Very good.

OpenStudy (anonymous):

Yes!!!! Thank you so much!! :) I truly do appreciate it :)

jimthompson5910 (jim_thompson5910):

you're welcome